OpenStudy (dls):

A sample of milk diluted with water has a density 1032 kg/m^3..if pure milk has a density of 1080 kg/m^3..find %of water by volume in milk!

OpenStudy (dls):

@shubhamsrg @Mashy

OpenStudy (dls):

kal test hai and this is the first question!!tell me something atleast :/!

OpenStudy (anonymous):

consider total volume to be X let y amt be milk.. then x-y amt is water..

OpenStudy (shubhamsrg):

final density = total mass/total volume

OpenStudy (shubhamsrg):

let milk mass be Mm, let water mass be Mw, then total mass M=Mw + Mm and total volume = x =>1032 = Mw + Mm /x also, 1080 = Mm/y and D = Mw/(x-y)

OpenStudy (shubhamsrg):

From last eqn, y=x- (Mw /D) putting that in eqn above that, Mm= ( x -(Mw/D) )*1080 putting this in 1st eqn, 1032 = (Mw/x) + (x- (Mw/D)*1080 No idea after this :P

OpenStudy (shubhamsrg):

We might well have to take same approximations at this point of time.